# Mapping tuberculosis fatalities by region and age group in South Korea: A dataset for targeted health policy optimization

### v1

@@ Abstract
In South Korea, age-disaggregated tuberculosis (TB) data at the district level are not publicly available due to privacy constraints, limiting fine-scale analyses of healthcare accessibility. To address this limitation, we present a high-resolution, district-level dataset on tuberculosis (TB) fatalities and hospital accessibility in South Korea, covering the years 2014 to 2022 across 228 districts. The dataset is constructed using a reconstruction method that infers age-disaggregated TB cases and fatalities at the district level by integrating province-level age-specific statistics with district-level spatial and demographic data, enabling analyses that account for both spatial heterogeneity and age structure. Building on an existing hospital allocation framework, we extend the objective function to an age-weighted formulation and apply it to the reconstructed dataset to minimize TB fatalities under different age-weighting schemes. We demonstrate that incorporating age structure can give rise to distinct optimized hospital allocation patterns, even when the total number of minimized fatalities is similar, revealing trade-offs between efficiency and demographic targeting. In addition, the dataset supports temporal analyses of TB burden, hospital availability, and demographic variation over time, and provides a testbed for spatial epidemiology and optimization studies that require high-resolution demographic and healthcare data.
Introduction
@@
Tuberculosis (TB) remains one of the leading causes of infectious disease mortality worldwide, with an estimated 1.3 million deaths annually [1, 2]. Despite long-term global efforts to control the disease, substantial regional disparities in TB incidence and outcomes persist [3, 4]. Within the Organization for Economic Co-operation and Development (OECD), South Korea presents a notable case: it consistently reports over 10,000 new TB cases annually and ranks second in TB incidence and fifth in TB-related mortality among OECD countries [5]. This combination of persistently high burden and comprehensive public health surveillance distinguishes Korea from most high-income countries with small numbers.
@@
South Korea maintains a fully digitized public health reporting system in which annual statistics on TB patients, fatalities, and healthcare infrastructure are collected and released at the administrative district level [6, 7], covering over 200 districts nationwide. In such data-rich settings, where disease surveillance is systematically curated and released with high spatial coverage, the integration of population and geographic information enables detailed analyses of spatial inequality and accessibility in urban systems [3, 4, 8, 9, 10]. South Korea provides a representative example of this environment, allowing us to examine hospital distribution and accessibility patterns in urban infrastructure using established spatial and optimization-based approaches [11, 12, 13, 3, 14].
More broadly, the spatial distribution of public infrastructure has been studied as a generic feature of urban systems, where facility density scales with population density [10, 13, 15, 12, 16]. In the context of healthcare, such spatial organization can influence not only average accessibility but also the robustness of service provision under uneven demand or localized constraints, as observed in other infrastructure systems [9, 17]. However, in practice, privacy constraints prevent most existing analyses from accessing disease data that are jointly disaggregated by age and administrative district at fine spatial resolution, limiting their ability to capture demographic heterogeneity in disease outcomes.
@@
Age is a critical determinant of TB disease severity [5, 18]. Older individuals are more likely to experience fatal outcomes and may face structural barriers to early diagnosis or continuous treatment [19, 20]. In South Korea, TB fatalities are heavily concentrated among individuals aged 70 years and older [Fig. 1(a)], whereas the general population is predominantly concentrated in middle-aged groups, indicating a pronounced mismatch between demographic structure and TB mortality risk that cannot be captured by age-agnostic models. However, as mentioned above, due to privacy regulations, Korean public health statistics do not provide TB data that are jointly disaggregated by age and higher-spatial-resolution administrative districts (e.g., specific age groups within specific districts) [6]. While province-level age disaggregation is available at lower spatial resolution, this limitation creates a structural data gap that hinders age-aware spatial analysis and equity assessments at the district level, where policy interventions are typically implemented [6].
@@
To address this gap, we reconstruct a district-level, age-disaggregated dataset of TB patients and fatalities spanning 2014 to 2022. Our method integrates publicly available province-level age distributions with district-level totals through an upscaling procedure, thereby introducing age resolution while preserving spatial fidelity. Because the reconstruction is based solely on aggregated statistics and does not involve individual-level records, it does not increase privacy risks beyond those present in the original data. The reconstructed dataset can also be used as a testbed for evaluating age-aware spatial analyses under realistic administrative constraints. The full reconstruction pipeline and source code are released alongside the dataset, enabling reproducibility and reuse in spatial epidemiology, healthcare accessibility, and infrastructure optimization studies that require both fine spatial resolution and demographic detail.
Furthermore, hospital accessibility plays a central role in disease outcomes, as effective treatment requires early diagnosis and repeated visits to secondary care facilities over extended periods [20, 19, 2]. Previous optimization studies have shown that increasing hospital density and accessibility is associated with reduced TB fatalities [11]. These approaches, however, typically assume uniform vulnerability across patients and do not incorporate age-specific fatality risk, despite strong empirical evidence that TB outcomes vary sharply with age.
Building on this line of work, we extend an existing hospital relocation optimization framework by embedding an age-weighted fatality minimization objective. By explicitly accounting for age-specific vulnerability, the model allows us to examine how demographic structure interacts with spatial allocation under fixed infrastructural constraints. We demonstrate that even when total minimized fatalities are comparable, incorporating age structure can lead to substantially different hospital allocation patterns across districts. By combining district-level data reconstruction with age-aware spatial optimization, this study provides a quantitative framework for analyzing how demographic heterogeneity influences healthcare accessibility and TB fatality outcomes in aging societies.
Background and Motivation
To optimize hospital density for minimizing TB fatalities, a previous study [11] formulated the TB fatalities $E$ as a function of hospital density while accounting for human mobility, based on a random-walk model with traps [21], as follows:
where the hospital density $\eta_{s}$ and the fatality rate  [^1] The case fatality rate is generally defined as the proportion of diagnosed patients who die from a disease, but this measure can be biased by undiagnosed cases and unresolved outcomes. For TB, estimates of the total number of cases are known to strongly correlate with the number of newly reported cases. Accordingly, in this study and in previous studies, the fatality rate is approximated by the ratio of TB-related deaths to newly reported TB cases in a given year. $\phi_{s}$ for district $s$ are defined as $\eta_{s}\equiv H_{s}/A_{s}$ and $\phi_{s}\equiv D_{s}/N_{s}$, where $H_{s}$ is the number of hospitals, $A_{s}$ is the district area, and $D_{s}$ and $N_{s}$ are the numbers of TB-related deaths and newly reported TB cases, respectively, in a given year. The hospital configuration across all districts is denoted by $\vec{\eta}=(\eta_{1},\eta_{2},\cdots,\eta_{s},\cdots)$. The characteristic hospital density $\tilde{\eta}_{s}$, which reflects the medical and infrastructural environment, can be empirically determined by equating the observed fatality rate ($\phi_{s}\equiv D_{s}/N_{s}$) to the modeled form $\phi_{s}=\exp(-\eta_{s}/\tilde{\eta}_{s})$, yielding $\tilde{\eta}_{s}=\eta_{s}/\log(N_{s}/D_{s})$, as reported in the previous study [11]. Preserving the total number of hospitals, the optimal hospital density of district $s$ is theoretically obtained as $\eta_{s}^{\rm(opt)}=\tilde{\eta}_{s}\log\left({\rho_{s}}/{(z\tilde{\eta}_{s})}\right)$, where $\rho_{s}={N_{s}}/{A_{s}}$ denotes the patient density and $z$ is a Lagrange multiplier. The previous study further demonstrated that the ratio ${\eta^{\rm(opt)}}/{\eta}$ between the optimal and current hospital densities strongly depends on the characteristic patient density ${\rho_{s}}/{\tilde{\eta}_{s}}$ [11].
@@
Figure 1: (a) Quartile plots of the fatality rate $\phi_{t}$ by age group $t$ across 16 provinces in 2022 (excluding Sejong-si due to the absence of reported fatalities). (b) Average age of the general population across the districts. Metropolitan districts tend to have younger populations, while non-metropolitan districts are characterized by older populations. (c) The distribution of average age displays a heterogeneous pattern in South Korea, indicating an overconcentration of younger individuals in metropolitan areas.
This modeling framework assumes that the fatality rate $\phi$ depends solely on the characteristics of each district and therefore does not capture the well-known dependence of TB fatality on patient age [1, 18]. This simplification is primarily due to data limitations: as of 2022, South Korea comprises 228 “si-gun-gu” districts (municipal government level, higher-resolution administrative units), which belong to 17 “si-do” provinces (regional local government level, lower-resolution units). Each province includes as few as 2 districts (Jeju Special Self-Governing Province) and as many as 31 districts (Gyeonggi-do, part of the Seoul metropolitan area). While TB data at the district level are more suitable for statistically robust analysis, age-disaggregated TB data are only available at the province level due to privacy concerns. Figure 1(a) illustrates the age-dependent TB fatality rates for the year 2022. Data provided by the Korea Disease Control and Prevention Agency (KDCA) [6] and Statistics Korea (KOSIS) [7] include the number $N_{t}$ of newly reported TB cases and the number $D_{t}$ of TB-related deaths for each age group $t$, aggregated at the province level. As shown in Fig. 1(a), the fatality rate $\phi_{t}=D_{t}/N_{t}$ exhibits substantial variation across age groups, with median values ranging from 0.0219 to 0.1690. Age groups under 40 years are not shown, as they exhibit negligible fatality rates.
In addition to age-related variability, the wide range of fatality rates observed within each age group suggests the presence of substantial spatial heterogeneity. While age-disaggregated TB data are not available at the district level, spatial differences can be indirectly inferred from the age composition of the general population. As shown in Fig. 1(b), metropolitan districts, including the capital city Seoul, tend to have younger average populations than non-metropolitan districts, resulting in a skewed age distribution across regions. This demographic contrast implies that the age composition of TB patients is also likely to vary spatially at the district level. These considerations motivate the need for TB data that jointly capture age structure and spatial resolution, which we address by constructing a higher-resolution, age-inclusive dataset and extending the fatality rate formulation accordingly.
@@ Methods
We utilize publicly available datasets reported annually from 2014 to 2022, provided by KDCA [6] and KOSIS [7]. This study requires age-disaggregated demographic and TB-related information at the district level, the higher-resolution administrative unit. Over this nine-year period, data have been collected for 17 provinces and 228 districts. In contrast to demographic information, such as the number $N$ of newly reported TB cases and the number $D$ of TB-related deaths by age group, non-demographic information, including the number $H$ of hospitals, is available at the district scale. Here, “hospital” refers to a secondary care hospital equipped to provide appropriate treatment for TB. The annual totals of each dataset are summarized in Table 1, and a subset of the data is visualized regionally in Fig. 2.
@@
Table 1: Total annual records of the TB empirical data. The number $N$ of newly reported TB cases and the number $D$ of TB-related deaths are available at the province level due to privacy concerns. The number HHof hospitals is provided at the district level. Here “hospital” refers to a secondary care hospital, which is equipped to provide proper treatment for TB.
Figure 2: Temporal patterns of the raw data for (a) the number $N$ of newly reported TB cases, (b) the number $D$ of TB-related deaths, and (c) the number $H$ of hospitals. The number of new TB cases and deaths is showing a declining trend. The number of secondary care hospitals is slightly increasing in the Gyeonggi-do region. The 16 provinces are represented by distinct symbols.
Figure 2 shows the temporal patterns of these data for each province separately. Seoul-si and Gyeonggi-do exhibit markedly higher values across all measured quantities than other provinces, reflecting their combined population of nearly half of South Korea’s total population (approximately 22 million out of 50 million). The number $N$ of newly reported TB cases shows a consistent downward trend across all provinces [Fig. 2(a)], and the number $D$ of TB-related deaths also decreases with small fluctuations [Fig. 2(b)], indicating steady nationwide progress toward TB control. The number $H$ of hospitals remains nearly unchanged in most provinces, except for Gyeonggi-do. As a major metropolitan region surrounding the capital city Seoul, Gyeonggi-do has continued to experience population growth driven by suburbanization. This population increase has been accompanied by a growing demand for infrastructure, including hospitals.
Figure 3: Regional and age-grouped tendency of (a) TB patients fraction $n$, (b) TB deaths fraction $d$, and, as a comparison, (c) the general population demographic fraction $g$ for the respective province in 2022. Both the proportion of new TB cases and TB-related deaths tend to increase with age. The 16 provinces are represented by distinct symbols.
@@
where $N_{i}=\sum_{t}N_{i,t}$ and $D_{i}=\sum_{t}D_{i,t}$. Ages are categorized into ten-year groupings. Similarly, the quantities $N_{t}$ and $D_{t}$, used to compute the age-dependent fatality rate $\phi_{t}$ shown in Fig. 1(a), are calculated as $N_{t}=\sum_{i}N_{i,t}$ and $D_{t}=\sum_{i}D_{i,t}$, respectively. The resulting age-group fractions are plotted in Fig. 3. The TB-related fractions $n$ and $d$, shown in Figs. 3(a) and 3(b), increase monotonically with age. Notably, age dependence is more pronounced for TB-related deaths than for newly reported cases. Moreover, TB deaths in Fig. 3(b) contrast with the age distribution of the general population shown in Fig. 3(c): although the population in South Korea is primarily concentrated in the 40–60 year age range, TB-related deaths are disproportionately concentrated among individuals aged 70 years and older. These observations indicate that the age distribution of TB fatalities deviates systematically from the underlying population structure, emphasizing the need to explicitly account for age-dependent effects in the subsequent analysis.
Upscaling TB demographic data to high spatial resolution
The raw demographic data related to TB described in the previous subsection are officially available in an age-disaggregated format only at the provincial level. To enable a more fine-grained analysis while maintaining statistical robustness, we develop a method to reconstruct TB case and fatality counts at the district level. Specifically, we assume that the age distributions observed at the provincial level [Figs. 3(a) and 3(b)] also apply to the districts within each province. Then, we can estimate the numbers of TB cases and deaths in district $s$, belonging to province $i$, for age group $t$ using Eq. (2) as
@@
where $N_{s}$ and $D_{s}$ denote the age-aggregated numbers of newly reported TB cases and TB-related deaths in district $s$, respectively, that are available in KDCA and KOSIS. The reconstructed data satisfy $\sum^{\prime}_{s}N_{s,t}=N_{i,t}$ and $\sum^{\prime}_{s}D_{s,t}=D_{i,t}$, where the primed summation runs over districts $s$ belonging to province $i$. In this reconstruction framework, the age-group fractions are assumed to be homogeneous within each province, such that $n_{s,t}=n_{i,t}$ for all districts $s\in i$.
@@
To implement the optimal distribution of hospitals based on the random walk model as in the previous study [11], the numbers of TB patients, TB-related deaths, and hospitals of each district are used. We extend this formulation to incorporate age effects using Eq. (3). In addition, as mentioned earlier and shown in Fig. 3(b), the number of TB-related deaths is extremely low (nearly zero) for age groups under 40. Therefore, we exclude age groups under 40, resulting in five age groups being considered: $t\in\{1,2,3,4,5\}$ (e.g., $t=1$ denotes the age group “40404949”). In summary, the subset of data that satisfies these requirements is extracted from the full set of districts and is presented in Table 2. The number of districts included in the analysis ranges from 113 to 143, out of a total of 228 districts.
Results
Modeling the TB fatality rate and objective function
@@ Results
Using upscaled TB data in Eq. (3), we aim to minimize the total number of TB-related deaths by reallocating hospitals across districts while accounting for age-specific effects. To this end, we reformulate the objective function in Eq. (1) by decomposing total fatalities into age-group components. Specifically, the original fatality rate $\phi_{s}$ is extended to $\phi_{s,t}\equiv D_{s,t}/N_{s,t}$, representing the fatality rate for age group $t$ in district $s$. As stated in Eq. (1), the fatality rate was modeled as $\phi_{s}(\eta_{s})=\exp(-\eta_{s}/\tilde{\eta}_{s})$, where $\eta_{s}$ denotes the areal density of hospitals and $\tilde{\eta}_{s}$ represents a district-specific characteristic density. In our age-disaggregated formulation, individuals in district $s$ access a shared set of $H_{s}$ hospitals regardless of age, implying that the hospital density remains $\eta_{s,t}=\eta_{s}$. However, hospital accessibility may vary across age groups (e.g., due to differences in mobility and activity), resulting in an age-dependent characteristic density $\tilde{\eta}_{s,t}$. Accordingly, the age-specific fatality rate is modeled as

### v2

@@ Abstract
In South Korea, age-disaggregated tuberculosis (TB) data at the district level are not publicly available due to privacy constraints, limiting fine-scale analyses of healthcare accessibility. To address this limitation, we present a high-resolution, district-level dataset on tuberculosis fatalities and hospital accessibility in South Korea, covering the years 2014 to 2022 across 228 districts. The dataset is constructed using a reconstruction method that infers age-disaggregated TB cases and fatalities at the district level by integrating province-level age-specific statistics with district-level spatial and demographic data, enabling analyses that account for both spatial heterogeneity and age structure. The reconstructed dataset supports temporal analyses of TB burden, hospital availability, and demographic variation over time, and provides a resource for spatial epidemiology and healthcare accessibility studies that require both fine spatial resolution and demographic detail.
Background & Summary
@@
Tuberculosis (TB) remains one of the leading causes of infectious disease mortality worldwide, with an estimated 1.3 million deaths annually [1, 2]. Despite long-term global efforts to control the disease, substantial regional disparities in TB incidence and outcomes persist [3, 4]. Within the Organization for Economic Co-operation and Development (OECD), South Korea presents a notable case: it consistently reports over 10,000 new TB cases annually and ranks second in TB incidence and fifth in TB-related mortality among OECD countries [5]. This combination of persistently high burden and comprehensive public health surveillance distinguishes Korea from most high-income countries with small case numbers.
@@
South Korea maintains a fully digitized public health reporting system in which annual statistics on TB patients, fatalities, and healthcare infrastructure are collected and released at the administrative district level [6, 7], covering more than 200 districts nationwide. In such data-rich settings, the integration of population and geographic information enables detailed analyses of spatial inequality and accessibility in urban systems [3, 4, 8, 9, 10]. More broadly, the spatial distribution of public infrastructure has been studied as a generic feature of urban systems, where facility density scales with population density [10, 13, 15, 12, 16]. In the context of healthcare, such spatial organization can influence both accessibility and robustness under uneven demand or localized constraints [9, 17].
@@
Age is a critical determinant of TB disease severity [5, 18]. Older individuals are more likely to experience fatal outcomes and may face structural barriers to early diagnosis or continuous treatment [19, 20]. However, due to privacy regulations, Korean public health statistics do not provide TB data that are jointly disaggregated by age and higher-spatial-resolution administrative districts [6]. While province-level age disaggregation is available at lower spatial resolution, this limitation creates a structural data gap that hinders age-aware spatial analysis and equity assessments at the district level.
@@
To address this gap, we reconstruct a district-level, age-disaggregated dataset of TB patients and fatalities spanning 2014 to 2022. Our method integrates publicly available province-level age distributions with district-level totals through an upscaling procedure, thereby introducing age resolution while preserving spatial fidelity. Because the reconstruction is based solely on aggregated statistics and does not involve individual-level records, it does not increase privacy risks beyond those present in the original data. The full reconstruction pipeline and source code are released alongside the dataset, enabling reproducibility and reuse in spatial epidemiology and healthcare accessibility studies.
Figure 1 summarizes the demographic motivation for the dataset. The TB fatality rate, defined as the number of TB deaths divided by the number of TB patients, increases strongly with age [Fig. 1(a)]. Since the average age of the general population varies substantially across districts [Fig. 1(b,c)], TB fatality outcomes are expected to differ across regions when age structure is taken into account. These observations motivate analyses of TB burden and healthcare accessibility using data that jointly capture age structure and spatial resolution.
@@
Figure 1: (a) Quartile plots of the TB fatality rate $\phi_{t}$ by age group $t$ across 16 provinces in 2022 (excluding Sejong-si due to the absence of reported fatalities). (b) Average age of the general population across districts. Metropolitan districts tend to have younger populations, while non-metropolitan districts are characterized by older populations. (c) Distribution of the district-level average age, showing demographic heterogeneity across South Korea.
@@ Methods
This section describes the data sources and reconstruction procedure used to generate the dataset. We describe how the district-level age-disaggregated TB data are constructed from the raw statistics provided by KDCA and KOSIS. The overall procedure is summarized in Fig. 2.
Figure 2: Flowchart of the data construction process. Starting from the raw source data (green), we reconstruct the district-level age-disaggregated quantities (pink). By construction, the reconstructed and raw quantities satisfy the relations $N_{s}=\sum_{t}N_{s,t}$ and $D_{s}=\sum_{t}D_{s,t}$. The dataset up to this stage corresponds to the records summarized in Table 1. For subsequent analyses, the dataset is further filtered using empirical criteria. The final dataset (yellow) contains districts $s$ that satisfy $H_{s}\geq 1$ and $N_{s,t},D_{s,t}\geq 1$ for all age groups $t$ of 40 years and older, corresponding to Table 2.
@@
We utilize publicly available datasets reported annually from 2014 to 2022, provided by KDCA [6] and KOSIS [7]. Over this nine-year period, data have been collected for 17 provinces and 228 districts. Demographic information, such as the number $N$ of newly reported TB cases and the number $D$ of TB-related deaths by age group, is available at the provincial level, whereas non-demographic information, including the number $H$ of hospitals, is available at the district level. As of 2022, South Korea comprises 228 “si-gun-gu” districts (municipal government level, higher-resolution administrative units), which belong to 17 “si-do” provinces (regional local government level, lower-resolution units). Each province includes as few as 2 districts (Jeju Special Self-Governing Province) and as many as 31 districts (Gyeonggi-do, part of the Seoul metropolitan area). Here, “hospital” refers to a secondary care hospital equipped to provide appropriate treatment for TB. The annual totals of these source datasets are summarized in Table 1. The province-level TB counts and district-level hospital counts form the basis for the reconstructed dataset.
@@
Table 1: Total annual records of the TB source data. Age-disaggregated counts of newly reported TB cases (NN) and TB-related deaths ($D$), covering all age groups, are available only at the province level due to privacy concerns. The number $H$ of hospitals is provided at the district level. Here “hospital” refers to a secondary care hospital equipped to provide proper treatment for TB.
Age distribution statistics
@@
where $N_{i}=\sum_{t}N_{i,t}$ and $D_{i}=\sum_{t}D_{i,t}$. Ages are categorized into ten-year groupings. Similarly, the quantities $N_{t}$ and $D_{t}$, used to compute the age-dependent fatality rate $\phi_{t}$ shown in Fig. 1(a), are calculated as $N_{t}=\sum_{i}N_{i,t}$ and $D_{t}=\sum_{i}D_{i,t}$, respectively.
Reconstruction of district-level age-disaggregated data
The raw demographic data related to TB are officially available in an age-disaggregated format only at the provincial level. To enable a more fine-grained analysis while maintaining statistical robustness, we reconstruct TB case and fatality counts at the district level. For a district $s$ belonging to province $i$, we assume that the age-group fractions are homogeneous within the province, such that $n_{s,t}=n_{i,t}$. Under this assumption, the numbers of TB cases and deaths in district $s\in i$ for age group $t$ are estimated as
@@
where $N_{s}$ and $D_{s}$ denote the age-aggregated numbers of newly reported TB cases and TB-related deaths in district $s$, respectively, that are available in KDCA and KOSIS. The reconstructed data satisfy $\sum^{\prime}_{s}N_{s,t}=N_{i,t}$ and $\sum^{\prime}_{s}D_{s,t}=D_{i,t}$, where the primed summation runs over districts in province $i$.
Data filtering and inclusion criteria
As shown in Fig. 1(a), we present TB fatality rates only for age groups aged 40 and above, since the rates for younger groups are negligible. Given the steep age dependence observed in the figure, it is also reasonable to expect that fatality rates below age 40 are negligible. Therefore, we focus on five age groups, $t\in\{1,2,3,4,5\}$ (with $t=1$ denoting the age group “40–49”). In addition, district-year records are retained only when they satisfy the conditions of having at least one newly reported TB case ($N\geq 1$), at least one TB-related death among individuals aged 40 years and older ($D\geq 1$), and at least one hospital ($H\geq 1$). The resulting annual summary of reconstructed district-level data is presented in Table 2.
@@
Data Records
The dataset supporting this study has been deposited in the Dryad Digital Repository and is available for peer review at the private link provided below. The released dataset contains annual TB statistics across administrative districts in South Korea from 2014 to 2022. It includes district-level counts of hospitals, TB patients, and fatalities, together with age-stratified proportions of new cases and deaths reported at the provincial level and the reconstructed district-level age-disaggregated quantities obtained from Eq. (2).
At the source-data level, the dataset covers 17 provinces and 228 districts (Table 1). The Dryad repository contains three province-level source files and two sets of district-level files. The province-level files provide annual TB fatalities, newly reported TB cases, and hospital counts for 17 provinces from 2014 to 2022, corresponding to Table 1. Province-level source files provide counts of fatalities and newly reported cases by 5-year age groups, whereas hospital counts are provided without age stratification.
At the district level, the repository provides one full reconstructed file for each year (sigungu_nd_age_YYYY.csv, 2014–2022) and one filtered analysis file for each year (sigungu_nd_age_filtered_YYYY.csv, 2014–2022). The full district-level files (Table 1) reproduce the province-level counts, except for Sejong-si, when aggregated to the provincial level. It contain district names, province names, hospital counts, district area, total newly reported TB cases, total TB-related fatalities, and reconstructed age-group proportions for new cases and fatalities. The filtered files (Table 2) are derived from the full district-level files and retain only districts with at least one newly reported TB case, at least one TB-related death, and at least one hospital, while restricting the age-resolved variables to age groups 40 years and older. At the reconstructed district-year level, the number of included districts ranges from 113 to 143 per year after applying the inclusion criteria summarized in Table 2.
In the full district-level files, the variables n0_9 to n80_ denote the age-group proportions of newly reported TB cases, and d0_9 to d80_ denote the age-group proportions of TB-related fatalities in Eq. (1), for 10-year age bins from 0–9 to 80+. The filtered district-level files include the corresponding variables only for age groups 40–49 through 80+, consistent with the analytical scope of the present study. District area is reported in square kilometers, and hospital variables correspond to the total number of hospitals recorded for each administrative district and year. A complete description of file contents, variable names, and age-group definitions is provided in the accompanying README and repository metadata, allowing users to distinguish source-level provincial counts from reconstructed district-level age-resolved quantities.
Technical Validation
Figure 3: Temporal patterns of the source data for (a) the number $N$ of newly reported TB cases, (b) the number $D$ of TB-related deaths, and (c) the number $H$ of hospitals. The 16 provinces with sufficient data are represented by distinct symbols.
Figure 4: Regional and age-grouped tendency of (a) TB patient fraction $n$, (b) TB death fraction $d$, and, for comparison, (c) the general population demographic fraction $g$ for each province in 2022. The 16 provinces are represented by distinct symbols.
Temporal consistency of source data
Figure 3 shows the temporal patterns of the source data for each province separately. Seoul-si and Gyeonggi-do exhibit markedly higher values across all measured quantities than other provinces, reflecting their large populations. The number of newly reported TB cases $N$ shows a consistent downward trend across provinces [Fig. 3(a)], and the number of TB-related deaths $D$ also decreases with small fluctuations [Fig. 3(b)]. The number of hospitals $H$ remains nearly unchanged in most provinces, except for Gyeonggi-do, where hospital counts increase modestly during the study period. These temporal patterns are preserved in the released dataset.
Age distribution verification
The age-group fractions defined in Eq. (1) are shown in Fig. 4. The TB-related fractions $n$ and $d$ increase monotonically with age, with a stronger age dependence for TB-related deaths than for newly reported cases [Fig. 4(a,b)]. For comparison, the age distribution of the general population is shown in Fig. 4(c). Although the population is concentrated in the 40–60 year age range, TB-related deaths (cases) are disproportionately concentrated among individuals aged 70 years and older. This comparison indicates that the increasing age pattern of TB-related quantities is not simply a reflection of the underlying population structure. Instead, it reveals a distinct age pattern specific to TB outcomes. These observations motivate the use of TB-related age fractions, rather than the general population age distribution, when constructing the age-disaggregated dataset. These age-specific patterns are also consistent with the decision to focus the filtered analysis dataset on individuals aged 40 years and older, for whom TB-related fatality is non-negligible.
Consistency of reconstructed counts
The reconstructed district-level data satisfy the aggregation constraints in Eq. (2) by construction. Summing $N_{s,t}$ and $D_{s,t}$ across districts within each province reproduces the province-level totals $N_{i,t}$ and $D_{i,t}$ reported in the original statistics. This ensures that the reconstructed district-level age-resolved dataset remains fully consistent with the officially reported province-level counts. In addition to the aggregate consistency described above, the released data files were organized so that the provenance of each variable can be traced directly to the underlying source or reconstruction step. Province-level files contain the original annual TB statistics by province, while district-level files store the reconstructed age-resolved quantities together with district-level totals of hospitals, TB cases, fatalities, and area. The filtered district-level files were generated deterministically from the full district-level files by applying the same inclusion criteria used in this paper.
Usage Notes
The reconstructed dataset is intended for comparative and exploratory studies of TB burden, healthcare accessibility, and demographic heterogeneity across South Korea. Because age-disaggregated district-level counts are inferred from province-level age fractions, the reconstructed quantities should be interpreted as estimates under the assumption of within-province homogeneity in age composition. Hospital accessibility in the dataset is represented through district-level hospital counts and areal hospital density. Such quantities provide a useful proxy for infrastructure availability, but they do not explicitly account for regional differences in hospital capacity, quality of care, or transportation infrastructure. The dataset may therefore be combined with other mobility or infrastructure data in future studies. In addition, the dataset can be used together with previously developed spatial allocation frameworks that analyze district-level hospital accessibility and TB fatalities [11].
Potential research application example: Healthcare facility optimization
As an illustrative application, the dataset can support studies on the spatial optimization of healthcare facilities. For example, the dataset can be combined with the hospital allocation framework proposed in Ref. [11]. To optimize hospital density for minimizing TB fatalities, they formulated the TB fatalities $E$ as a function of hospital density while accounting for human mobility, based on a random-walk model with traps [21], as follows:
where the hospital density $\eta_{s}$ and the fatality rate $\phi_{s}$ for district $s$ are defined as $\eta_{s}\equiv H_{s}/A_{s}$ and $\phi_{s}\equiv D_{s}/N_{s}$, where $H_{s}$ is the number of hospitals, $A_{s}$ is the district area, and $D_{s}$ and $N_{s}$ are the numbers of TB-related deaths and newly reported TB cases, respectively, in a given year. The hospital configuration across all districts is denoted by $\vec{\eta}=(\eta_{1},\eta_{2},\cdots,\eta_{s},\cdots)$. The characteristic hospital density $\tilde{\eta}_{s}$, which reflects the medical and infrastructural environment, can be empirically determined by equating the observed fatality rate ($\phi_{s}\equiv D_{s}/N_{s}$) to the modeled form $\phi_{s}=\exp(-\eta_{s}/\tilde{\eta}_{s})$, yielding $\tilde{\eta}_{s}=\eta_{s}/\log(N_{s}/D_{s})$, as reported in the previous study [11]. Preserving the total number of hospitals, the optimal hospital density of district $s$ is theoretically obtained as $\eta_{s}^{\rm(opt)}=\tilde{\eta}_{s}\log\left({\rho_{s}}/{(z\tilde{\eta}_{s})}\right)$, where $\rho_{s}={N_{s}}/{A_{s}}$ denotes the patient density and $z$ is a Lagrange multiplier. The previous study further demonstrated that the ratio ${\eta^{\rm(opt)}}/{\eta}$ between the optimal and current hospital densities strongly depends on the characteristic patient density ${\rho_{s}}/{\tilde{\eta}_{s}}$ [11]. This modeling framework assumes that the fatality rate $\phi$ depends solely on the characteristics of each district and therefore does not capture the well-known dependence of TB fatality on patient age [1, 18]. This simplification is primarily due to data limitations. While TB data at the district level are more suitable for statistically robust analysis, age-disaggregated TB data are only available at the province level due to privacy concerns.
Figure 5: (a) Illustrative example of how the released dataset can be used in spatial healthcare allocation models. District-level values of the difference ${(\eta^{\rm(opt)}_{\rm age}-\eta^{\rm(opt)}_{\times})}/{\eta}$ across South Korea are visualized. The inset shows a magnified view of Seoul. Uncolored districts indicate no data satisfying the criteria. (b-c) Comparison of the difference ${(\eta^{\rm(opt)}_{\rm age}-\eta^{\rm(opt)}_{\times})}/{\eta}$ for $\rho_{s,5}$ (b) and $\rho_{s,5}/\tilde{\eta}_{s,5}$ (c). No distinct pattern observed along $\rho_{s,5}$, while a steady, roughly linear increase is seen along $\rho_{s,5}/\tilde{\eta}_{s,5}$ despite the log-scaled x-axis. This figure is provided solely as an illustrative example of dataset usage.
@@
Using age-resolved district-level TB data in Eq. (2), it is possible to minimize the total number of TB-related deaths by reallocating hospitals across districts while accounting for age-specific effects. Reformulating the objective function in Eq. (3) by decomposing total fatalities into age-group components, the original fatality rate $\phi_{s}$ is extended to $\phi_{s,t}\equiv D_{s,t}/N_{s,t}$, representing the fatality rate for age group $t$ in district $s$. As stated in Eq. (3), the fatality rate was modeled as $\phi_{s}(\eta_{s})=\exp(-\eta_{s}/\tilde{\eta}_{s})$, where $\eta_{s}$ denotes the areal density of hospitals and $\tilde{\eta}_{s}$ represents a district-specific characteristic density. In the age-disaggregated formulation, individuals in district $s$ access a shared set of $H_{s}$ hospitals regardless of age, implying that the hospital density remains $\eta_{s,t}=\eta_{s}$. However, hospital accessibility may vary across age groups (e.g., due to differences in mobility and activity), resulting in an age-dependent characteristic density $\tilde{\eta}_{s,t}$. Accordingly, the age-specific fatality rate is modeled as